OpenStudy (anonymous):

If you place a bag 3/4 full of sugar on one pan of a balance scale. On the other pan you place an identical bag 2/3 full of sugar along with a 1/5 pound weight. If the scale balances, how many pounds does a full bag of sugar weight? Explain

OpenStudy (anonymous):

3w/4 = 2w/3 +1/5 w/12 = 1/5 w = 12/5 A full bag of sugar weighs 2.4 lbs.

OpenStudy (anonymous):

let x be required no. of pounds 3/4 x = 2/3 x + 1/5 1/12 x = 1/5 x = 12/5 full bag weighs 2.4 pounds